OpenStudy (anonymous):

Determine the derivative of 3x^2 at x=6 using the definition of derivative

OpenStudy (anonymous):

36

OpenStudy (unklerhaukus):

\[\frac{\text d}{\text dx} y(x)=\frac{\text d}{\text dx} 3x^2=6x\] \[\frac{\text d}{\text dx} y(6)=6(6)=36\]

OpenStudy (anonymous):

Thanks!!!